OpenStudy (anonymous):

find the value of that that makes the trinomial a perfect square: x^2+3x+c

OpenStudy (anonymous):

9/4

OpenStudy (anonymous):

9/4

OpenStudy (amistre64):

we take the 'x' term and divide it in half; then square it (3/2)^2 = 9/4

OpenStudy (amistre64):

\(x^2 + (b)x + (\frac{b}{2}*\frac{b}{2})\) is the format